The Happy Savings Bank offers a nominal interest rate of 6% compounded quarterly, whereas the People's Credit Union offers a nominal interest rate of 6.2% compounded semi annually. Suppose you want to open a new bank account and you have narrowed your choices to one of these institutions.

What is the Effective Interest Rate for Happy Savings Bank?
What is the Effective Interest Rate for People's Credit Union?
Where should you open your account?

Answer by LinnW(1048)

For Happy, (1 + 6%/4)^4 = ( 1 + 1.5%)^4 = 1.015^4 = 1.061336..
so their effective rate is 6.1336%.
For People's , (1 + 6.2%/2)^2 = (1 + 3.1%)^2 = 1.031^2 = 1.06296..
so their effective rate is 6.296%
Based upon what we know, we will earn more at People's so
if everything else is the same, we should open the account at People's.